Please help me solve this equation:
cubed root of (x+2) = sixth root of (9x+10).

Answer by user_dude2008(1862)

cubed root of (x+2) = sixth root of (9x+10).
(x+2)^(1/3) = (9x+10)^(1/6)
(x+2)^(1/3) = ((9x+10)^(1/2))^(1/3)
x+2 = (9x+10)^(1/2)
(x+2)^2 = 9x+10
x^2+4x+4 = 9x+10
x^2+4x+4-9x-10 = 0
x^2-5x-6 = 0
(x-6)(x+1)=0
x=6 or x=-1 <-------- Answers